A pool has a perimeter of 176 ft, and a length that is three time the width. What's the pool's area?
Answer by macston(5194)

.
W=width; L=length=3W; P=perimeter
.
P=2(L+W)
176ft=2(3W+W)
88ft=4W
22ft=W The width is 22 feet.
L=3W=3(22ft)=66ft The length is 66 feet.
.
Area=LW
Area=(66ft)(22ft)=1452 sq ft
ANSWER: The area of the pool is 1452 square feet.